9b^3 + 7b - 5*(9 b^3 - 2b) =
Answer by lenny460(766)

Simplify:
9b^3 + 7b - 5(9b^3 - 2b)
-5 * 9b^3 = -45b^3
-5 * -2b = 10b
Therefore:
9b^3 + 7b - 45b^3 + 10b
Combine like terms:
9b^3 - 45b^3 = -36b^3
7b + 10b = 17b
The Answer:
17b - 36b^3


obuong3@aol.com